# Clinicians’ Voice: Fundamental Considerations for XAI in Healthcare

### v2

@@ 1 Introduction
Rapid developments in artificial intelligence (AI) call for more transparency of AI-based technologies to tackle concerns related to trustworthiness, responsibility, and accountability. This has led to the emergence of the field of Explainable AI (XAI). The goal of XAI is making AI models more transparent and understandable to address these concerns and aid the implementation of AI-based technologies. Especially in high-stakes situations, like those in healthcare, AI-based systems could enhance critical decision-making processes (e.g., Eke and Shuib, 2024; Alowais et al., 2023). Ideally, AI models provide a certain prediction (e.g., the probability of a skin lesion being cancerous) which supports the clinicians in their decision process. However, implementation in the medical field – beyond regulatory requirements these technologies have to meet – has been slow partly due to lacking transparency (Eke and Shuib, 2024; He et al., 2019). Although user studies, evaluations, and multidisciplinary research are generally advocated for within XAI, only a small fraction of publications evaluate their proposed solutions with users. For example, Saarela and Podgorelec (2024) reviewed 512 articles, where almost 60% did not do any evaluation at all, and only 8% performed qualitative evaluation involving domain experts. Similarly, Nauta et al. (2023) report that only one in five papers, of over 300 papers they reviewed, performed evaluations with users. Arguably, XAI approaches to medical imaging as used in radiology may be most frequently implemented, as shown by the number of publications with XAI applications in this sub-field alone (e.g., Borys et al., 2023; Gupta and Seeja, 2024). For tabular data, however, there are less XAI applications, implementations and user tests. Therefore, our work focuses on tabular data structures like electronic health records. Specifically, in this research, we conduct semi-structured interviews with clinicians from the Netherlands. Our aim is to elicit clinicians’ thoughts on XAI, and by that, uncover compelling perspectives that may give rise to new research avenues and considerations in the design of XAI methods for healthcare and especially for tabular data structures. For example, feature importance scores are often praised for their simplicity, however our research suggests that they may become unbelievable due to exactly this property. Furthermore, our interviewees also raised concerns about taking responsibility for decisions, even if there is sufficient trust in the AI.
@@
Overall, our interview study provides compelling statements and unique angles that put existing research into a new light and may give rise to new research avenues. We also find support for previous research, including that users emphasize the importance of receiving information about the background of the model in addition to its predictions and explanations. Our results further highlight the importance of multidisciplinary projects to ensure seamless introduction of (X)AI-based tools into clinical practice, and the significance of training clinicians on AI before implementing AI-based tools in their workflow.
@@ 2 Background
Healthcare is a high-stakes domain with vast opportunities for successful (X)AI implementations. This is also reflected in the increasing number of publications on AI-based clinical decision support tools (DSTs) that incorporate explainability. For example, Thoral et al. (2021) trained an AI-model to predict readmission or death within seven days after discharge from the ICU and explain predictions using feature importance scores (Lundberg and Lee, 2017). Both Abdullah et al. (2021) and Antoniadi et al. (2021) provide reviews of explainable machine learning approaches in healthcare including examples of publications on medical AI tools with explainability. Markus et al. (2021) focus on how explainability can be evaluated and provide a step-by-step guide to choose the most appropriate XAI method. Amann et al. (2020) take a multidisciplinary approach to determine the need for explainability in healthcare from technological, legal, medical, patient, and ethical perspectives. They conclude that there is an apparent need for explainability and highlight the importance of the interaction between healthcare professionals and AI-based DSTs.
@@
Our study differs from this previous research in several ways. First, we do not build our work around a specific case or AI-tool. In doing so, we eliminate the factor of expertise with the specific device or case from the study design. Besides, while we also consider this type of research important, we think it is just as relevant to take a step back and use a holistic approach for XAI in healthcare. We aim to elicit and analyze perspectives of clinicians that may be universal to a range of AI products and fields within healthcare. Second, in line with that, we interview clinicians from different fields within healthcare to record general data. Third, our study is exploratory in nature rather than empirically driven. That is, we do not test for an effect of XAI methods on, for example, trust or understanding. Qualitative interview studies are known to be exploratory in nature and generalizing findings may be difficult. Nonetheless, they pose a useful tool to gather in-depth understanding that can enrich results of quantitative studies and to uncover unique perspectives and compelling angles for further research (Tenny et al., 2022; Dunwoodie et al., 2023). The goal of our research is to do just that, as well as to identify clinicians’ requirements for successful implementation of AI-tools in clinical practice with the aim of improving decision making in healthcare.
@@ 3 Methodology
We conducted 11 semi-structured interviews with open-ended questions with clinicians trained and working in the Netherlands. In this section we discuss recruitment strategy, interview procedure, materials, and limitations of our design. The supplementary materials for the interviews are available in the appendix. The study was approved by the ethical body of our institution [the details will be filled after the review process in order to preserve the anonymity of our submission]. Interviewees were informed of the goal and procedure of the interviews and signed an informed consent form prior to their participation.
@@ 3.1 Participants and recruitment
The main inclusion criterion for participants was successful completion of their medical study in the Netherlands. The sample included clinicians from different fields within medicine (general practice, (pediatric) oncology, radiology, general health psychology, obstetrics and gynecology, surgery) and with years of experience in their fields ranging from two to 30 years. Participants were recruited through networking and word of mouth. None of the interviewees was personally known to the interviewer before the interview was conducted.
@@ 3.2 Procedure
Potential interviewees were contacted via email, which stated the reason for reaching out, the person that had provided the contact details, practical information such as the expected time, and the informed consent form. If interviewees responded positively, an exchange followed to arrange a meeting. Interviews were scheduled for 30 minutes, and the majority of them took place online as participants come from across the Netherlands. During the interview, the interviewer confirmed with participants that they agree to have the interview audio-recorded. Interviewees could ask questions or withdraw from the interview at any moment, and were explicitly given the opportunity to add any information they wished to mention or ask questions at the end of the interview. All interviews were conducted by the same interviewer to avoid potential differences due to interviewing style.
@@ 3.3 Materials
An interview scheme with guiding questions was developed to provide structure and guidance to the interviews (see Table 1). Broadly, aspects that we looked at include (i) tasks that clinicians (would) find (X)AI helpful for, (ii) desirable characteristics of XAI methods/explanations, (iii) hopes as well as concerns and fears they face regarding the implementation of (X)AI in their workflow, and (iv) their perception of the role of trust in the implementation and adoption of (X)AI.
@@
At the start of each interview, the interviewer provided context for the interview after introductions, similar to what participants had read in the informed consent form beforehand. Then, the first set of questions was aimed at inquiring about the interviewee’s background, including their specialization and experience. The following questions were aimed at understanding the current use of AI in their field and healthcare in general. Besides, we also asked interviewees to share about research in (X)AI for healthcare that they are aware of. Then, we encouraged them to come up with use cases of AI in their field which they would find helpful in their job. Following up on that, we asked them for reasons why these do not yet exist, and why implementation of AI in healthcare in general is slow. The next set of questions was focused on uncovering interviewee’s attitude towards AI in healthcare. This entailed asking specifically for their positive and negative feelings towards using AI in healthcare, and if they could describe their hopes and concerns. In this part of the interview, we also asked them about the role they believe trust plays in the implementation and adoption of (X)AI in practice. Then, for the next set of questions, the interviewer explained three different types of explanations using a simple illustrative example for illustration (see Appendix A for the exact material and information given to participants) and asked interviewees to explain their thoughts considering each of these explanation types. The explanations that we chose were (1) feature importance scores, (2) counterfactual explanations, and (3) a glass box model.
@@
Counterfactual explanations show the minimal change required in the input feature space to change the prediction of a sample to, for example, a different class. In the literature, the sample and its corresponding prediction that is to be explained is coined factual, whereas the changed data point leading to the desired prediction is called counterfactual. Research from the social and cognitive sciences shows that counterfactual explanations mimic the way humans give explanations themselves, and hence are expected to be good explanations in XAI (Byrne, 2019; Miller, 2019), which has spurred many researchers to propose algorithms generating these explanations (e.g., Russell, 2019; Ustun et al., 2019; Kanamori et al., 2020, 2021; Karimi et al., 2020, 2021; Mothilal et al., 2020; Mahajan et al., 2019; Maragno et al., 2022).
@@
Interviewees received a concise but detailed information sheet as supplementary material before the interview, which included the illustrative example and the three explanations, complemented with information about what they show and how to interpret them. We include the material that interviewees received in Appendix A. Besides their opinion on these specific explanation types, we included a question on what characteristics (e.g., visual, interactive) they would like to see in an AI system, or what type of information they would like to be shown. Then, at the end of the interview, interviewees were asked if they would like to share any other considerations with us.
@@ 3.5 Limitations
The nature of our study bears some limitations we should consider when interpreting the results. First, interviews are known to bear threats to validity and reliability (Alshenqeeti, 2014). As a response to this, we chose to have the same interviewer conduct all interviews. Further, interviews were audio-recorded and then transcribed for the analysis, by that avoiding misinformation through note-taking or incorrect understanding by the interviewer. Furthermore, during the interviews, time was taken to clarify any unclarities to avoid misunderstanding and misinterpretation of the responses. Second, as an explorative interview study, the sample size is limited and we did not perform any quantitative analyses of the results and should be careful with making generalizations. Third, the interviews were scheduled to last 30 minutes, which may not be considered long enough to discuss different aspects of XAI. As this is an exploratory study and clinicians are known to have limited time, we were constrained by this duration. Lastly, we reached out to candidates via email, and in some cases a common connection made the introduction. Generally, it can be expected that individuals interested in XAI are more likely to positively respond to an interview request. Hence, we may only have interviewed individuals with strong (positive or negative) opinions on the topic, by that missing out on the broader group of clinicians.
@@ 4 Results
In this section we discuss patterns identified through our thematic analysis of the interviews and put them into perspective using direct quotes and examples from interviewees.
@@ 4.1 Current use of AI in healthcare
Interviewees generally talked about AI-based tools in healthcare in three different categories, namely tools related to (1) business operations, (2) care processes, and (3) medical device regulations. The first category of AI systems is relatively universal to any larger business, not only hospitals or care facilities, and covers aspects like finance, forecasting, and so on. The second category relates to AI systems that enhance the care process by, for instance, making it more efficient. A typical example that several interviewees provided was the planning of chemotherapy. The last category of AI systems in healthcare concerns AI systems that fall under the medical device regulation. Simply put, a medical device is any instrument intended to be used to diagnose, treat, monitor, or prevent diseases “and does not achieve its primary intended action by pharmacological, immunological, or metabolic means” (International Medical Device Regulators Forum (2013), IMDRF, p.7). AI models fall under the category Software as a Medical Device, which includes software that is intended to be used for any of the aforementioned purposes without being part of a hardware medical device (International Medical Device Regulators Forum (2013), IMDRF). In layman’s terms, decision support tool is more commonly used to describe AI as a medical device. Here, an AI model is trained to make a prediction based on patient data and hence, supports clinicians in their decisions.
When asked to what extent AI models are implemented in medical care in their field specifically, interviewees unanimously said that there is very little AI (as DSTs) integrated in their workflows. Most interviewees pointed out that a significant amount of budget is invested in research and development of medical AI models. Although many of these efforts are being tested, very few of them actually reach the bedside. Arguably, the most popular field for AI applications within healthcare is radiology, with the number of publications on AI in radiology being “sky high” (interview 4). However, even though there are more than 200 products in AI for radiology available on the European market [^1] https://radiology.healthairegister.com, not many of those are integrated in clinical practice yet. Furthermore, one interviewee pointed out that most of these products perform specific subtasks, such as segmentation or quantification tasks, instead of predicting a concrete medical outcome; thus, only indirectly supporting clinicians.
However, some interviewees also mentioned that other forms of AI models, such as those for people planning and scheduling, are integrated everywhere, and considering AI in that sense “everybody is really already using AI a lot”, but “not […] as a medical device yet” (interview 10). Note that this falls into the second category of AI models previously mentioned.
@@ 4.2 What tasks should AI be used for?
Even though some interviewees held the opinion that AI for administrative tasks is already widely used, others said that they would want to see AI be more used for administrative and logistics tasks. One interviewee said that “80% of our task is administration” (interview 6), hence if AI could be used for this type of work, clinicians could focus more on their patients. Two interviewees mentioned a type of large language model for medical purposes, for example to help them write summary reports which is a tedious task taking up a significant amount of time.
Regarding AI as a clinical DST, four interviewees mentioned predicting complications as a useful application. For example, when performing surgery, they would like to know upfront how large the chance is to run into a certain complication. And even beyond that, one interviewee said they “would like to know upfront what can go wrong and how can we prevent that” (interview 5). Hence, AI could serve not only as a warning system but also as a guide on how to navigate certain situations. To give a concrete example, one interviewee from obstetrics and gynecology explained that when a woman is in labor, they monitor both the mother and the baby, meaning they collect a large dataset. Having an AI model trained on this data that continuously predicts the chance of a birth ending in a C-section would save them not only time and resources, but would also reduce the risks for women in labor. Another example comes from oncology, where it would be beneficial to predict the likelihood of complications during surgery but also the probability of remission following treatment.
Two interviewees mentioned that monitoring or medical surveillance would be helpful, especially in cases where patients are not continuously at the hospital. With the recent technological advancements, it may be possible to provide a wide range of patients with wearables to monitor their health. In this case, (X)AI-based tools could be used to predict time of admission to the hospital or make recommendations on further treatment, which could then be evaluated in consultations and check-ups with the clinician.

### v3

@@ 1 Introduction
Rapid developments in artificial intelligence (AI) call for more transparency of AI-based technologies to tackle concerns related to trustworthiness, responsibility, and accountability. This has led to the emergence of the field of Explainable AI (XAI). The goal of XAI is making AI models more transparent and understandable to address these concerns and aid the implementation of AI-based technologies. Especially in high-stakes situations, like those in healthcare, AI-based systems could enhance critical decision-making processes (e.g., Eke and Shuib, 2024; Alowais et al., 2023). Ideally, AI models provide a certain prediction (e.g., the probability of a skin lesion being cancerous) which supports the clinicians in their decision process. However, implementation in the medical field – beyond regulatory requirements these technologies have to meet – has been slow partly due to lacking transparency (Eke and Shuib, 2024; He et al., 2019). Although user studies, evaluations, and multidisciplinary research are generally advocated for within XAI, only a small fraction of publications evaluate their proposed solutions with users. For example, Saarela and Podgorelec (2024) reviewed 512 articles, where almost 60% did not do any evaluation at all, and only 8% performed qualitative evaluation involving domain experts. Similarly, Nauta et al. (2023) report that only 20% of over 300 papers they reviewed performed evaluations with users. Arguably, XAI approaches to medical imaging as used in radiology may be most frequently implemented, as shown by the number of publications with XAI applications in this sub-field alone (e.g., Borys et al., 2023; Gupta and Seeja, 2024). For tabular data, however, there are less XAI applications, implementations and user tests. Therefore, our work focuses on tabular data structures like electronic health records. Specifically, in this research, we conduct semi-structured interviews with clinicians from the Netherlands. Our aim is to elicit clinicians’ thoughts on XAI and, by that, uncover compelling perspectives that may give rise to new research avenues and considerations in the design of XAI methods for healthcare and especially for tabular data structures. For example, feature importance scores are often praised for their simplicity, however our research suggests that they may become unbelievable due to exactly this property. Furthermore, our interviewees also raised concerns about taking responsibility for decisions, even if there is sufficient trust in the AI.
Generalizable Insights about Machine Learning in the Context of Healthcare
@@
Overall, our interview study provides compelling statements and unique angles that put existing research into a new light and could inspire new research. We also find support for previous research, including that users emphasize the importance of receiving information about the background of the model in addition to its predictions and explanations. Our results further highlight the importance of multidisciplinary projects to ensure seamless introduction of (X)AI-based tools into clinical practice, and the significance of training clinicians on AI before implementing AI-based tools in their workflow.
Studies like this one are often overlooked in machine learning research albeit their importance for developing holistic AI solutions for healthcare. Generalizing findings and claims about interpretability of AI models requires keeping clinicians in the loop, and multidisciplinary collaboration at all stages of the design and development process is key to successful implementation and bringing AI to the bedside.
@@ 2 Background
Healthcare is a high-stakes domain with vast opportunities for successful (X)AI implementations. This is also reflected in the increasing number of publications on AI-based clinical decision support tools (DSTs) that incorporate explainability. For example, Thoral et al. (2021) trained an AI-model to predict readmission or death within seven days after discharge from the ICU and explain predictions using feature importance scores (Lundberg and Lee, 2017). Abdullah et al. (2021) and Antoniadi et al. (2021) provide reviews of explainable machine learning approaches in healthcare, and Markus et al. (2021) provide a step-by-step guide to choose the most appropriate XAI method. Amann et al. (2020) take a multidisciplinary approach to establish the need for explainability in healthcare from technological, legal, medical, patient, and ethical perspectives.
While it is generally expected that complementing AI systems with explanations will increase task performance and trust (e.g., Senoner et al., 2024; Lim et al., 2009), there is also research suggesting that explanations complementing AI outputs may lead to over-reliance. For example, van der Waa et al. (2021) and Jabbour et al. (2023) show that AI-based DSTs with explanations do not help participants identify incorrect predictions. Alarmingly, Kaur et al. (2020) showed that even data scientists tend to over-rely on AI systems if explanations are provided. These results emphasize the need for a careful design process considering contextual requirements and the interplay of humans and technology.
@@
Our study differs from this previous research in several ways. First, we do not build our work around a specific case or AI-tool. We aim to elicit and analyze perspectives of clinicians that may be universal to a range of AI products and fields within healthcare. Second, we interview clinicians from different fields within healthcare to record general data. Third, our study is exploratory in nature rather than empirically driven. That is, we do not test for an effect of XAI methods on, for example, trust or understanding. Qualitative interview studies are known to be exploratory in nature and generalizing findings may be difficult. Nonetheless, they pose a useful tool to gather in-depth understanding that can enrich results of quantitative studies and to uncover unique perspectives and angles for further research (Tenny et al., 2022; Dunwoodie et al., 2023).
@@ 3 Methodology
We conducted 11 semi-structured interviews with open-ended questions with clinicians trained and working in the Netherlands. In this section we discuss recruitment strategy, interview procedure, materials, and limitations of our design. The supplementary materials for the interviews are available in the appendix. The study was approved by the ethical body of our institution, the Economics and Business Ethics Committee of the University of Amsterdam, The Netherlands. Interviewees were informed of the goal and procedure of the interviews and signed an informed consent form prior to their participation.
@@ 3.1 Participants and recruitment
The sample consisted of six women and five men, with years of experience in their fields ranging from two to 30 years. The main inclusion criterion for participants was successful completion of their medical study in the Netherlands. The sample included clinicians from different fields within medicine (general practice, (pediatric) oncology, radiology, general health psychology, obstetrics and gynecology, surgery). Participants were recruited through networking and word of mouth. None of the interviewees was personally known to the interviewer before the interview was conducted.
@@ 3.2 Procedure
Potential interviewees were contacted via email, which stated the reason for reaching out, the person that had provided the contact details, practical information such as the expected time, and the informed consent form. If interviewees responded positively, an exchange followed to arrange a meeting. Interviews were scheduled for 30 minutes, and the majority of them took place online. During the interview, the interviewer confirmed with participants that they agree to have the interview audio-recorded. Interviewees could ask questions or withdraw from the interview at any moment, and were explicitly given the opportunity to add any information at the end of the interview. All interviews were conducted by the same interviewer to avoid potential differences due to interviewing style.
@@ 3.3 Materials
An interview scheme with guiding questions was developed to provide structure and guidance to the interviews (see Table 1). Broadly, aspects that we looked at include (i) tasks that clinicians (would) find (X)AI helpful for, (ii) desirable characteristics of XAI methods/explanations, (iii) hopes, concerns, and fears they face regarding the implementation of (X)AI in their workflow, and (iv) their perception of the role of trust in the implementation and adoption of (X)AI.
@@
At the start of each interview, the interviewer provided context for the interview after introductions, similar to what participants had read in the informed consent form beforehand. Then, the first set of questions was aimed at inquiring about the interviewee’s background, including their specialization and experience. The following questions were aimed at understanding the current use of AI in their field and healthcare in general. Besides, we also asked interviewees to share about research in (X)AI for healthcare that they are aware of. Then, we encouraged them to come up with use cases of AI in their field which they would find helpful in their job. Following up, we asked them for reasons why these do not yet exist, and why implementation of AI in healthcare in general is slow. The next set of questions was focused on uncovering interviewee’s attitude towards AI in healthcare. This entailed asking specifically for their positive and negative feelings towards using AI in healthcare, and if they could describe their hopes and concerns. In this part of the interview, we also asked them about the role they believe trust plays in the implementation and adoption of (X)AI in practice. Then, for the next set of questions, the interviewer explained three different types of explanations using a simple illustrative example for illustration (see Appendix B for the exact material and information given to participants) and asked interviewees to explain their thoughts considering each of these explanation types. The explanations that we chose were (1) feature importance scores, (2) counterfactual explanations, and (3) a glass-box model.
@@
Counterfactual explanations show the minimal change required in the input feature space to change the prediction of a sample to a different class. In the literature, the sample and its corresponding prediction that is to be explained is coined factual, whereas the changed data point leading to the desired prediction is called counterfactual. Research from the social and cognitive sciences shows that counterfactual explanations mimic the way humans give explanations, and hence are expected to be good explanations in XAI (Byrne, 2019; Miller, 2019). This has spurred many researchers to propose algorithms generating these explanations (e.g., Russell, 2019; Ustun et al., 2019; Kanamori et al., 2020, 2021; Karimi et al., 2020, 2021; Mothilal et al., 2020; Mahajan et al., 2019; Maragno et al., 2022).
@@
Interviewees received a concise but detailed information sheet as supplementary material before the interview, which included the illustrative example and the three explanations, complemented with information about what they show and how to interpret them. We include the material that interviewees received in Appendix B. Besides their opinion on these specific explanation types, we included a question on what characteristics (e.g., visual, interactive) they would like to see in an AI system, or what type of information they would like to be shown. Then, at the end of the interview, interviewees were asked if they would like to share any other considerations with us.
@@ 3.5 Limitations
The nature of our study bears some limitations we should consider when interpreting the results. First, interviews are known to bear threats to validity and reliability (Alshenqeeti, 2014). As a response to this, we chose to have the same interviewer conduct all interviews. Further, interviews were audio-recorded and then transcribed for the analysis, by that avoiding misinformation through note-taking or incorrect understanding by the interviewer. Furthermore, during the interviews, time was taken to clarify any unclarities to avoid misunderstanding and misinterpretation of the responses. Second, as an explorative interview study, the sample size is limited and we did not perform any quantitative analyses of the results and should be careful with making generalizations. Third, the interviews were scheduled to last 30 minutes, which may not be considered long enough to discuss different aspects of XAI. As this is an exploratory study and clinicians are known to have limited time, we were constrained by this duration. Lastly, we reached out to candidates via email, and in some cases a common connection made the introduction. Generally, it can be expected that individuals interested in XAI are more likely to positively respond to an interview request. Hence, we may only have interviewed individuals with strong (positive or negative) opinions on the topic.
@@ 4 Results
We next discuss patterns identified through our thematic analysis of the interviews and put them into perspective using direct quotes and examples from interviewees. While reserving our main discussion for Section 5, we discuss in this section the relevant literature directly linking to interviewee’s responses.
@@ 4.1 Current use of AI in healthcare
Interviewees generally talked about AI-based tools in healthcare in three different categories, namely tools related to (1) business operations, (2) care processes, and (3) medical device regulations. The first category of AI systems is relatively universal to any larger business, not only hospitals or care facilities, and covers aspects like finance, forecasting, and so on. The second category relates to AI systems that enhance the care process by, for instance, making it more efficient (e.g., the planning of chemotherapy). The last category of AI systems in healthcare concerns AI systems that fall under the medical device regulation. Simply put, a medical device is any instrument intended to be used to diagnose, treat, monitor, or prevent diseases “and does not achieve its primary intended action by pharmacological, immunological, or metabolic means” (International Medical Device Regulators Forum (2013), IMDRF, p.7). AI models fall under the category Software as a Medical Device and are commonly referred to as decision support tools, meaning that the AI system’s output is intended to support clinicians in their decision process.
When asked to what extent AI models are implemented in medical care in their field specifically, interviewees unanimously said that there is very little AI (as DSTs) integrated in their workflows. Most interviewees (7/10) [^1] Due to the nature of semi-structured interviews not every question was posed to each interviewee. Therefore, the number of respondents may differ per question and we indicate it accordingly. pointed out that a significant amount of budget is invested in research and development of medical AI models. Although many of these efforts are being tested, very few of them actually reach the bedside. Arguably, the most popular field for AI applications within healthcare is radiology, with the number of publications on AI in radiology being “sky high” (interview 4). However, even though there are more than 200 products in AI for radiology available on the European market [^2] https://radiology.healthairegister.com, not many of those are integrated in clinical practice yet. Furthermore, one interviewee pointed out that most of these products perform specific subtasks, such as segmentation or quantification tasks, instead of predicting a concrete medical outcome; thus, only indirectly supporting clinicians.
However, some interviewees (3/9) also mentioned that other forms of AI models, such as those for people planning and scheduling, are integrated everywhere, and considering AI in that sense “everybody is really already using AI a lot”, but “not […] as a medical device yet” (interview 10). Note that this falls into the second category of AI models previously mentioned.
@@ 4.2 What tasks should AI be used for?
Even though some interviewees held the opinion that AI for administrative tasks is already widely used, others (4/9) said that they would want to see AI be more used for administrative and logistics tasks. One interviewee said that “80% of our task is administration” (interview 6), hence if AI could be used for this type of work, clinicians could focus more on their patients. Two interviewees mentioned a type of large language model for medical purposes, for example to help them write summary reports which is a tedious task taking up a significant amount of time.
Regarding AI as a clinical DST, four out of nine interviewees mentioned predicting complications as a useful application. For example, when performing surgery, they would like to know upfront how large the chance is to run into a certain complication. And even beyond that, one interviewee said they “would like to know upfront what can go wrong and how can we prevent that” (interview 5). Hence, AI could serve not only as a warning system but also as a guide on how to navigate certain situations. To give a concrete example, one interviewee from obstetrics and gynecology explained that when a woman is in labor, they monitor both the mother and the baby, meaning they collect a large dataset. Having an AI model trained on this data that continuously predicts the chance of a birth ending in a C-section would save them not only time and resources, but would also reduce the risks for women in labor. Another example comes from oncology, where it would be beneficial to predict the likelihood of complications during surgery but also the probability of remission following treatment.
Two out of nine interviewees mentioned that monitoring or medical surveillance would be helpful, especially in cases where patients are not continuously at the hospital. With the recent technological advancements, it may be possible to provide a wide range of patients with wearables to monitor their health. In this case, (X)AI-based tools could be used to predict time of admission to the hospital or make recommendations on further treatment, which could then be evaluated in consultations and check-ups with the clinician.